### v1

# Enhancing key rates of QKD protocol by Coincidence Detection

Abstract
In theory, quantum key distribution (QKD) provides unconditional security; however, its practical implementations are susceptible to exploitable vulnerabilities. This investigation tackles the constraints in practical QKD implementations using weak coherent pulses. We improve on the conventional approach of using decoy pulses by integrating it with the coincidence detection (CD) protocol. Additionally, we introduce an easy-to-implement algorithm to compute asymptotic key rates for the protocol. Furthermore, we have carried out an experimental implementation of the protocol, where we demonstrate that monitoring coincidences in the decoy state protocol leads to enhanced key rates under realistic experimental conditions.
I Introduction
[…]
QKD offers information-theoretic security, allowing two authenticated parties to exchange information securely through a secret and random key. The first QKD protocol was proposed in 1984 Bennett1984 followed by many others E91 ; BBM92 ; DPSK ; COW ; MDI ; Lucamarini2018 . The principles of quantum mechanics guarantee security without relying on any assumptions regarding the adversary’s capabilities Lo1999 ; Shor2000 ; Mayers2001 . While QKD promises unconditional security, its practical implementations are fraught with numerous vulnerabilities that a potential eavesdropper can exploit Huttner1994 ; Brassard2000 ; Vakhitov2001 ; Makarov2005 ; Gisin2006 ; Lydersen2010 ; Jain2011 ; Biswas2021 .
The imperfect sources provide one such loophole. In most implementations of prepare and measure protocols, weak coherent pulses (WCPs) are used instead of a single photon source (SPS). These weak coherent pulses follow Poissonian statistics; hence, there is a non-zero probability of multiphoton pulses. An adversary can exploit the presence of multiple photons in a pulse to perform a photon number splitting (PNS) attack to gain information on the shared key.
To address the exposed vulnerabilities in the practical implementation of QKD protocols, we can either transition to a more robust protocol or apply security patches for a known attack. Security patching involves monitoring various parameters of our system and assessing potential information leaks. By accounting for imperfections in the source, we can estimate the information leakage caused by pulses containing more than one photon. Security proofs that take into account device imperfections were introduced in Refs. Lutkenhaus2000 ; Gottesman ; Inamori2007 . Additionally, the authors in Ref. Hwang2003 proposed employing decoy states as a defence against the PNS attack. By integrating the work of Gottesman and Hwang2003 , the framework for the decoy-state protocol was built Lo2005 ; Wang2005 ; Ma2005 , which is one of the most widely implemented QKD protocols.
Coincidence monitoring is another alternative strategy to mitigate a PNS attack. The expected and observed coincidences at the receiver’s end are compared for a source emitting weak coherent pulses with known photon statistics and a well-characterized channel. Earlier studies felix_faint_2001 ; biswas_quantum_2022 ; Lim:17 have proposed monitoring the coincidences to limit the eavesdropper’s information on the key, while some of them also suggest using these coincidences to enhance the key rate. However, Lutkenhaus_2002 ; PhysRevResearch.5.023065 suggests that an eavesdropper can perform a sophisticated PNS attack, where she can mimic the detector statistics.
In the present study, we propose combining decoy state protocol with coincidence detection to circumvent a large class of such sophisticated PNS attacks. Following the common convention, we refer to the sender as Alice, the receiver as Bob, and the adversary as Eve. We monitor the coincidences not only for the signal but also for the decoy states. Since Eve does not have the information on whether a signal or decoy is sent, it is not possible for Eve to practically mimic the statistics at the receiver’s end. This is done in addition to the standard decoy state protocol, hence adding an extra layer of security. Therefore, even if Eve manages to accurately reproduce the statistics on average at Bob’s side using most strategies, this will result in a discrepancy in the yield between the signal and the decoy pulses. By incorporating two-fold coincidences, we can achieve higher key rates while enhancing security against sophisticated PNS attacks facilitated by decoy states. The integration of decoy states and coincidence detections helps us to strategically leverage their capabilities to our advantage. We will call this decoy state coincidence detection (DSCD) protocol.
If the obtained coincidence distribution indicates that a PNS attack is not carried out, we incorporate the contributions from two-photon gain ($Q_{2}$) and error ($e_{2}$) into the asymptotic key rate(refer to Lim:17 for the derivation of key rates in terms of $Q_{2}$ and $e_{2}$). Further, three or higher photon gains and error terms cannot be incorporated if we require security against the unambiguous state discrimination (USD) attack. Having additional contributions from the two-photon statistics makes it necessary to have a technique to compute tight bounds on the key rates that can be obtained in the protocol.
Computing the key rate of a general QKD protocol is a difficult problem. This is because the process of computing key rates for the protocol requires solving a nonlinear optimization problem involving many variables, for which tight lower bounds can often not be obtained using analytic techniques. Note that while it may be possible (on a case-by-case basis) to compute lower bounds on the key rates by making clever approximations on the individual parameters of the optimization problem, such as done by Ma et al. in their seminal paper Ma2005 , these approximations often lead to pessimistically underestimated key rates.
Obtaining tight bounds on the key rates based on the generalized decoy state protocol has been of interest, and several novel techniques decoy-SUH ; decoy-SZC ; decoy-ZNTC ; decoy-AHNPT ; decoy-BGKE ; decoy-CML ; decoy-GLL have been proposed that successfully compute good lower bounds on the rates. The key idea often used in the literature is to recast the decoy state protocol in terms of an entanglement-based protocol and then perform the security analysis of such a protocol. These techniques are based on sophisticated results in convex optimization and are incredibly powerful tools when computing key rates of a general QKD protocol.
Here, we present a simple-to-implement method to obtain (in principle) tight bounds on the key rates for our proposed protocol as well as the decoy state protocol. Our method diverges from the approaches discussed above and relies on casting the optimization problem for computing the key rate into a sequence of converging polynomial optimization problems, each of which can be efficiently solved using well-known numerical techniques. Our method is generally easy to implement and simple to code.
In principle, while it is possible to compute arbitrarily tight bounds on the key rate, achieving greater precision demands a higher amount of computation time and resources. However, in practice, our method produces reasonably tight key rates within a few minutes on a personal computer with a standard configuration. The proposed protocol and the method to calculate the key rates are then employed in a free-space QKD implementation. We employ our technique to demonstrate that our proposed protocol outperforms the decoy state protocol when the coincidence monitoring allows us to certify the absence of the PNS attack.
This research article is organized as follows. In Sec. II, we introduce the protocol, and in Section III, we discuss our method to compute the secret key rate for the protocol. Sec. IV describes the experimental setup and data analysis method. We discuss the results of our experiment and our implementation of SDP for computing key rates in Sec. V and end with some concluding remarks in Sec. VI.
Figure 1: Standard Scheme for Quantum Communication in free space with existing classical systems
II Protocol
The implementation of the BB84 protocol is susceptible to PNS attacks, where Eve exploits occasional multi-photon pulses generated by Alice’s source. The conventional countermeasure to the PNS attack, namely the decoy state protocol, involves modifications at Alice’s end by sending occasional decoy pulses alongside signal pulses. An alternative approach to combat the PNS attack is to monitor the coincidences at Bob’s end, as suggested by the works of Lim et al. Lim:17 and Biswas et al. biswas_quantum_2022 . The protocol proposed in this paper implements modifications at both Alice’s and Bob’s ends. Decoy states are sent from Alice’s end, and Bob’s end is equipped to monitor coincidence detections, effectively addressing PNS attacks and significantly enhancing key rates. For clarity, the protocol is systematically stated below:
Protocol (Decoy State Coincidence Detection Protocol (DSCD)).
Here, we outline a general decoy state protocol, including coincidences monitoring. Alice transmits a signal with mean photon number $\nu_{0}$ accompanied by $K$ random decoy pulses with mean photon numbers $\nu_{1},\nu_{2},\cdots\nu_{K}$. Given a sequence of $n$ pulses, each labelled with an index $i$, the following steps are repeated for each pulse.
  1. 1.
    State preparation (Alice’s lab): Alice generates random numbers $\mathcal{D}_{i}\in\{0,1,\cdots K\}$, $\mathcal{X}_{i}\in\{0,1\}$ and $\mathcal{A}_{i}\in\{0,1\}$. Where,
    1. (a)
      $\mathcal{D}_{i}$ determines whether a signal ($\nu_{0}$) or a decoy-state ($\nu_{1},\nu_{2},\cdots\nu_{K}$) is transmitted. If $\mathcal{D}=d$ then mean photon number is $\nu_{d}$
    2. (b)
      $\mathcal{X}_{i}$ determines the basis in which pulse $i$ is encoded. If $\mathcal{X}_{i}=0(1)$, she encodes in Standard (Hadamard) basis.
    3. (c)
      $\mathcal{A}_{i}$ is the bit value encoded for pulse $i$.
  2. 2.
    State transmission: Alice sends the prepared state to Bob through an authenticated quantum channel.
  3. 3.
    State measurement (Bob’s lab): Bob lets the pulse pass through a beam splitter, having two ports $\mathfrak{B}^{(0)}$ and $\mathfrak{B}^{(1)}$. At wing, $\mathfrak{B}^{(0)}$ ($\mathfrak{B}^{(1)}$) Bob performs measurement in the Standard (Hadamard) basis. Let $\mathcal{B}_{i}^{(j)}$ denote the measured bit value for the pulse $i$ in port $j$.
  4. 4.
    Alice and Bob continue to prepare and measure, incrementing the value of $i$ to $i+1$ until $i=n$. Once $i=n$, they proceed to the next step.
  5. 5.
    Alice publicly discloses the value of her bases $(\mathcal{X}_{1},\mathcal{X}_{2},\cdots,\mathcal{X}_{n})$.
  6. 6.
    Let $\mathcal{Y}_{i}$ denote Bob’s basis choice and $\mathcal{B}_{i}$ is the measured bit value. No measurement outcome is denoted as $\perp$.
    1. (a)
      If $\mathcal{B}_{i}^{(j\oplus 1)}=\perp$ and $\mathcal{B}_{i}^{(j)}\in\{0,1\}$: $\mathcal{Y}_{i}=j$ , $\mathcal{B}_{i}=\mathcal{B}_{i}^{(j)}$.
    2. (b)
      If $\mathcal{B}_{i}^{(0)},\mathcal{B}_{i}^{(1)}\in\{0,1\}$, then Bob sets $\mathcal{Y}_{i}=x$ and $\mathcal{B}_{i}=\mathcal{B}_{i}^{(x)}$ if $\mathcal{X}_{i}=x$.
    3. (c)
      Else Bob sets $\mathcal{Y}_{i}$ randomly and $\mathcal{B}_{i}=\perp$.
  7. 7.
    Bob publicly disclose the value of the bases $(\mathcal{Y}_{1},\mathcal{Y}_{2},\cdots\mathcal{Y}_{n})$ respectively. Alice and Bob discard the rounds in which $\mathcal{X}_{i}\neq\mathcal{Y}_{i}$ i.e. they only focus on the rounds from the set $\mathfrak{h}:=\{i\in\{1,2,\cdots n\}:\mathcal{X}_{i}=\mathcal{Y}_{i}\}$.
  8. 8.
    Alice also discloses the decoy states used in each round $(\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n})$.
  9. 9.
    Alice and Bob perform error correction and privacy amplification as in standard BB84 protocol. From any randomly chosen subset $\mathfrak{g}\subset\mathfrak{h}$, with $|\mathfrak{g}|\ll|\mathfrak{h}|$ they estimate Gains (probability that a signal sent by Alice is received by Bob):
    $Q_{\nu_{d}}$ ::=:= |$\frac{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d,\mathcal{B}_{i}\neq\perp\}|}{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d\}|}$ (1)
    and QBERs (the probability that $\mathcal{A}_{i}\neq\mathcal{B}_{i}$ whenever $\mathcal{X}_{i}=\mathcal{Y}_{i}$ ):
    $E_{\nu_{d}}$ ::=:= |$\frac{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d,\mathcal{A}_{i}\neq\mathcal{B}_{i}\}|}{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d\}|}$ (2)
The standard decoy state protocol is typically executed similarly to protocol Protocol, utilizing one beam splitter and two wings $\mathfrak{D}^{(0)}$ and $\mathfrak{D}^{(1)}$. $\mathfrak{D}^{(0)}$ performs measurements in the standard basis, while $\mathfrak{D}^{(1)}$ conducts measurements in the Hadamard basis. The main difference lies in the monitoring of coincidences, thus not affecting the ease of experimental implementation of the protocol. In Section IV, we present an experimental realization of this protocol.
In the standard decoy state protocol, the key rate can be computed in terms of the single photon gain $Q_{1}$ and error rate $e_{1}$ (see appendix A for definitions) as:
$R\geq\frac{1}{2}\left\{-Q_{\mu}H_{\text{bin}}(E_{\mu})f(E_{\mu})+Q_{1}\left(1-\Phi\left(2e_{1}-1\right)\right)\right\}.$ (3)
Here, $H_{\mathrm{bin}}(x):=x\log_{2}(x)-(1-x)\log_{2}(x)$ is the binary entropy and
$\Phi(x)$ ::=:= $H_{\mathrm{bin}}\left(\frac{1}{2}+\frac{x}{2}\right).$ (4)
In their work Lim et. al. Lim:17 , have argued that, with coincidence monitoring, keys can also be extracted from $2$ photon pulses. They argue that if Eve performs the Unambiguous State Discrimination (USD) attack, then it fails with certainty on two-photon pulses, whereas it may succeed with a non-zero probability for $3$ or more photon pulses. In our protocol, the photon number statistics of the two-photon pulses are monitored via coincidence detections. Therefore, we can be certain if the two-photon statistics are unaltered or not. If these statistics are altered, then we assume the PNS attack and rates can be computed by (3). However, if the coincidences detected show the expected statistics of coincidences, then we can safely assume that the PNS attack is not carried out, and the key rate of Protocol Protocol can be calculated using the following bound (given in Lim:17 ):
$R\geq\frac{1}{2}\left\{-Q_{\mu}H_{\text{bin}}(E_{\mu})f(E_{\mu})+Q_{1}\left(1-\Phi\left(2e_{1}-1\right)\right)+Q_{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\right\}.$ (5)
Here, $Q_{1}=Y_{1}\mu$ and $Q_{2}$ and $e_{2}$ are the two-photon gains and error rates (see appendix A), respectively. A similar argument was also presented in Biswas2021 . When computing the secure key rates for the protocol, the gains $Q_{1}$ and $Q_{2}$, as well as the error rates $e_{1}$ and $e_{2}$, are unknown, and one must account for the most pessimistic case, where the adversary is the most powerful. However, in either reference, the method to compute secure key rates was not addressed. In the next section, we provide a simple-to-implement numerical method that computes secure key rates for the protocol.
The standard decoy state protocol usually outperforms the BB84 protocol in key rate due to improved estimation of single-photon statistics ($Q_{1}$ and $e_{1}$). Decoy state statistics aid in obtaining better estimates for both signal and decoy states, resulting in higher key rates. The same argument also applies in our protocol, where having the decoy states, in principle, also allows for better estimates for the single and two-photon statistics $Q_{1},Q_{2},e_{1}$ and $e_{2}$.
Note that, as pointed out in felix_faint_2001 ; biswas_quantum_2022 ; Lim:17 , the PNS attack itself can be detected by simply measuring the coincidences even in the absence of decoy states. However, there are modifications to the PNS attack that cannot be detected simply by monitoring coincidences. One such example is an attack in which Eve only allows for one photon to pass and then sends photons herself to match the ideal coincidence detection. However, such an attack is easily prevented by using decoy states. This is because the adversary does not fundamentally know whether the state sent is a decoy or not; therefore, she will not be able to match the coincidences correctly when Alice and Bob condition on the rounds in which the signal/decoy was sent.
However, the strongest form of the modifications in the PNS attack will still not be detected using our protocol. In this attack, Eve injects a photon only if she stores one. However, we argue that such an attack is not practical for Eve to implement given the current technology. For Eve to execute such an attack, she must first determine that she has stored a photon, generate a replacement photon, and then send it to Bob. This process, especially with current technology, is likely to create a detectable time delay, which Bob can identify. Therefore, ideally, Bob should utilize the time lag information to discard rounds with significant time discrepancies.
III Computing key-rate
To compute the key rate, we must minimize the key rate over all the strategies that achieve the observed experimental statistics. In other words, the key rate can be obtained by solving the following optimization problem:
$r:=\min$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$ (6)
$\mathrm{s.t.}$ $\forall k:\quad Y_{k},e_{k}\in[0,1]$
$\forall d:\quad Q_{\nu_{d}}e^{\nu_{d}}=\sum_{k=0}^{\infty}Y_{k}\frac{\nu_{d}^{k}}{k!}$
$\forall d:\quad E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}=\sum_{k=0}^{\infty}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}$
The reason for performing such an optimization is straightforward to understand. The objective function of the optimization problem concerns the key rate $R$ (refer to Eqn (5)). This key rate relies on the $k$ photon yields $Y_{k}$ and $k$ photon error rates $e_{k}$, which are not directly known from experimental statistics. However, what is known are the overall gain $Q_{\mu}$ and the QBER ($E_{\mu}$). The constraints of the optimization problem are related to the observed Gain and QBER for both the signal and decoy states. The fundamental idea is that the experimentally observed QBER and Gain should originate from certain, albeit unknown, fixed values of Yields and error rates.
Given that no assumptions can be made about the Yields $Y_{k}\in[0,1]$ and error rates $e_{k}\in[0,1]$, we acknowledge the possibility of a potential adversary to manipulate them freely. The adversary is, in principle, unrestricted in choosing any values for these parameters, provided that the chosen values align with the observed experimental statistics. This fact is captured using the constraints in the optimization problem, limiting the choice available to the adversary. To account for the strategy that is most advantageous to the adversary within these constraints, we minimize the rate. In the literature, Ma et al. Ma2005 conducted the first comprehensive exploration of lower bounds on the asymptotic key rates for the decoy state protocol. For the decoy state protocol without any coincidence detection, the optimization problem for computing the key rate differs from (6) only in the objective function. Specifically, it does not include contributions from the two-photon terms $Y_{2}$ and $e_{2}$ and is given by:
$Y_{1}\mu(1-\Phi(2e_{1}-1))=Y_{1}\mu H_{\mathrm{bin}}(e_{1}).$ (7)
Therefore, the rates obtained in our protocol cannot fall below those of the standard decoy state protocol. In our experiments, a significant increase in the key rate is observed, as discussed in Section V. For the case with only one decoy, traditional analytical methods compute lower bounds on the key rate by first establishing bounds on $Y_{1}$ and $e_{1}$:
$Y_{0}\leq Y_{0}^{U}$ ::=:= $\frac{E_{\nu}Q_{\nu}e^{\nu}}{e_{0}},$ (8)
$Y_{1}\geq Y^{L}_{1}$ ::=:= $\frac{\mu}{\mu\nu-\nu^{2}}\left(Q_{\nu}e^{\nu}-Q_{\mu}e^{\mu}\left(\frac{\nu^{2}}{\mu^{2}}\right)-Y_{0}^{U}\left(\frac{\mu^{2}-\nu^{2}}{\mu^{2}}\right)\right),$ (9)
$e_{1}\leq e_{1}^{U}$ ::=:= $\frac{E_{\nu}Q_{\nu}e^{\nu}}{Y_{1}^{L}\nu}$ (10)
Exploiting the monotonicity properties of binary entropy, specifically that $H_{\mathrm{bin}}(x)$ is increasing for all $x\in[0,1/2]$, and the fact that $(1-H_{\mathrm{bin}}(x))>0$, a lower bound on the secure key rate is given by:
$R\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\mathrm{bin}}(E_{\mu})+Y_{1}^{L}\mu e^{-\mu}\left(1-H_{\mathrm{bin}}(e_{1}^{U})\right)\Big{\}}.$ (11)
An immediate approach to computing lower bounds on the key rate for the protocol with coincidence detection involves employing similar analytic techniques to establish upper and lower bounds, denoted as $Y_{2}^{L}$ and $e_{2}^{U}$, on $Y_{2}$ and $e_{2}$ respectively, that are compatible with the constraints. However, this process is tedious, and tight bounds may not be easily found. While numerical methods can be employed to find such bounds (one approach to obtain such bounds is shown in Appendix D), the results obtained are not sufficiently tight, as demonstrated in (refer to Figure-2(a). In the next section, we introduce a simple-to-implement method to iteratively derive bounds on the key rate that converge to the asymptotic key rate from below.
III.1 Solving the optimization problem
Two primary challenges emerge when addressing optimization problem (6). First, the problem is non-convex due to the non-convex nature of both the constraints and the objective function in (6). General optimization problems are notoriously difficult to solve due to their abstract nature unless they fit specific classes of convex optimization problems. As a result, the optimization problem cannot be solved directly using conventional techniques. Second, optimization problem (6) involves infinitely many free parameters, namely, $Y_{k}$ and $e_{k}$. The presence of these infinitely many parameters adds further complexity to the problem.
The challenge of dealing with infinitely many variables can be addressed straightforwardly by recognizing that the contribution of $Y_{k}$ and $e_{k}$ when $k\gg 1$ is negligible to the sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ and $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}e_{k}$. Consequently, we relax the constraints of (6) by truncating the infinite sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ to involve sums over a finite number of variables. This comes, however, with a small penalty that depends on the number of terms kept in the sum. We formally implement this truncation in Appendix C.1.
We begin by discussing our general technique for addressing certain simple non-polynomial optimization problems, drawing inspiration from BhRC2023 ; bhavsar2023improvements . This method involves deriving lower bounds for optimization problems through the partitioning of the parameter space into smaller sub-spaces. Formally, this partition (or grid), denoted as $\mathcal{P}$, divides the space into multiple sub-domains $\{\mathcal{C}_{i}\}_{i}$. For each sub-domain $\mathcal{C}_{i}$, we formulate a new optimization problem (or sub-problem) by constraining the parameters to that sub-domain. Each sub-problem can be lower-bounded by a polynomial optimization problem using simple methods such as Taylor’s theorem, with only a minimal loss of tightness. Obtaining lower bounds on the resulting sub-problems is achieved by computing the Semidefinite Programming (SDP) relaxations of the polynomial optimization problem (in practice, this is done using software tools like NCPOL2SDPA NCPOL2SDPA ). A reliable lower bound on the optimization problem is then determined by identifying the lowest computed value (see Lemma 2, Appendix C.2). Detailed information on partitioning and lower-bounding the sub-problems generated by any given partition is formally done in Appendix C.3.
The solutions to the optimization problems can be improved by refining the partition so that the sub-domains have smaller dimensions. Typically, if the partitioning is done over the parameter space of $P$ independent parameters, then enhancing tightness by a factor of $\beta$ would require computing $\beta^{P}$ times more sub-problems. Consequently, this technique is particularly applicable and effective when partitioning is applied to a relatively smaller number of parameters.
Now let us return to the optimization problem at hand. There are only two non-polynomial terms $\Phi(2e_{1}-1)$ and $\Phi((2e_{2}-1)^{2})$ in the optimization problem. Furthermore, only two parameters, $e_{1}$ and $e_{2}$, contribute to the non-polynomial terms in the problem. As $e_{1}$ and $e_{2}$ are both constrained within the range of $[0,1]$, we construct a partition $\mathcal{P}$ of the set $[0,1]\times[0,1]$ , generating rectangular sub-domains $\mathcal{C}_{i}$. Leveraging the properties of the function $\Phi(x)$, we find the constants $\xi^{\max}_{1,i}$ and $\xi^{\max}_{2,i}$ that (tightly) lower bound the functions $\Phi(2e_{1}-1)$ and $\Phi((2e_{2}-1)^{2})$ respectively in the sub-domain $\mathcal{C}_{i}$ [^1] To minimize the objective function, a lower-bound on $-\Phi((2e_{k}-1)^{k})$ is needed. (See lemma 3, Appendix  C.3). We then use these tight bounds to lower-bound the objective function as follows
$\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}(1-\Phi((2e_{k}-1)^{k}))\geq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}\cdot(1-\xi^{\max}_{k,i}).$ (12)
The loss in tightness when bounding the objective functions by such constants $\xi^{\max}_{1,i}$ and $\xi^{\max}_{2,i}$ depend upon the dimensions of the domain $\mathcal{C}_{i}$ as shown in Lemma 3, Appendix C.3. Thus, the loss of tightness can be made arbitrarily small by refining the partition. However, making a finer partition comes at the expense of higher computation time, as more and more sub-problems need to be numerically solved.
All the results lead to the following final result which forms the basis of the algorithm for computing key-rate $r$ (as defined in equation (6)). We state it here informally as:
[…]
In essence, this result implies that the lower bounds can be determined with arbitrarily high precision by solving multiple polynomial optimization problems and then computing the minimum of all the results obtained. We give a formal statement of Theorem 1 along with its proof in Appendix C.3.
Assuming a random background with errors occurring equally likely, we explicitly set the background error rate as $e_{0}=1/2$. To account for the impact of finite statistics on the experimental values $Q_{\mu}$, $Q_{\nu}$, $E_{\mu}$, and $E_{\nu}$, it is imperative to consider uncertainties in their measurement. In Appendix E, we modify Theorem 1 to accommodate these statistical uncertainties in the computation of rates. It is important to note that the key rates reported here are key rates in the asymptotic limit. However, this analysis only serves to acknowledge the uncertainty associated with experimentally observed values, and the computed rates are still asymptotic key rates compatible with experimentally obtained statistics with relevant uncertainties in observed values. The computation of finite-round statistics can, in principle, be performed using theoretical tools such as the Generalized Entropy Accumulation Theorem GEAT ; GEAT-security . Such analysis would require a more sophisticated analysis and is reserved for future work.
IV Experimental Method
This section discusses the experimental methods for implementing our secure QKD protocol. The setup comprises three stages: state preparation, transmission and state measurement.
Figure 2: Schematic of the experimental setup. It includes both the optics and electronic components. LD: Laser Diodes, HWP, Half Wave Plate, BS: Beam Splitter, PBS: Polarising Beam Splitter, NDF: Neutral Density Filter, VOA: Variable optical Attenuator, DM: Dichroic Mirror, L: lens, MMF: Multi-Mode Fibre, SPCM: Single Photon Counting Module, DAQ: Data Acquisition system
IV.1 State Preparation: Alice
Alice uses the polarisation degree of freedom of weak coherent pulses to encode the quantum state to be sent. Weak coherent pulses are generated using laser diodes and attenuated using neutral density filters. Our experimental setup Fig. 2 comprises four laser diodes operating at a wavelength of $808$ nm (Thorlabs $L808P010$). An in-house designed laser driving circuit driven by a Field Programmable Gated Array (FPGA) activates the laser diodes, producing pulses at a frequency of $5$ MHz, each with an optical pulse width of $1$ ns. All four lasers are then optically engineered to generate the four polarisations: horizontal (H), vertical (V), diagonal (D) and anti-diagonal (A), using polarizing beam splitters (PBS) and half-wave plates (HWP). The combination of PBS and HWP acts as a polariser as well as aids in the attenuation of the pulse. The FPGA also ensures that the four lasers are randomly triggered, thereby ensuring that the four polarisation states are randomly generated. We have characterized all four sources to account for various side channels arising from discrepancies in various degrees of freedom such as pulse width, wavelength, and power Biswas2021 . After combining all the polarizations on the beam splitter BS, the signal is passed through one fixed and one variable neutral density filter, which is used to attenuate the beam further. The variable NDF is rotated to vary the intensity of the signal, measured in terms of mean photon number, characterised by using the method proposed in sharma_2024 .
IV.2 Transmission: The Channel
The signal transmission to Bob takes place through a free space channel. The communication framework within the Thaltej campus of the Physical Research Laboratory (PRL) in Ahmedabad, Gujarat, India, encompasses two neighbouring buildings. Both the transmitter and receiver stations are situated in adjacent rooms on the rooftop of the first building. A reflector is placed on the second building to direct the signal to Bob. We have conducted our experiment in the natural environment rather than a controlled laboratory setting, deliberately incorporating variables such as atmospheric interference and night lighting. This approach ensures a more comprehensive and realistic assessment, enhancing the credibility and applicability of our findings.
IV.3 State Measurement: Bob
Upon reaching Bob, the signal undergoes a decoding process involving projective measurements onto the four polarisation states. The choice of basis for the measurement is achieved by using a beam splitter (BS). The presence of an HWP @$22.5$ in one of the paths designates the diagonal basis, while the other path represents the rectilinear basis. The signals are detected using single photon counting modules (Excelitas-SPCM-AQRH-14), integrated with a high-performance data acquisition card to record the timestamps of detections across various polarisations. Given the inherent nature of our source as a WCP, a non-zero probability of coincidences exists during the detection process. After a sufficient exchange of quantum states, Alice and Bob proceed to classical postprocessing.
IV.4 Postprocessing
Alice publicly discloses the choice of her bases. Bob encounters single detections as well as coincidences. In the first case, Bob keeps those detections where his choice of basis matched with that of Alice. In the latter case, Bob is only interested in two-fold coincidences, specifically those which occur in opposite bases. In these situations, Bob chooses the measurement in the basis which is compatible with Alice’s choice. These steps are as described in Steps 5-8 of the protocol described in Sec. II. These refined bits constitute the sifted key. A part of this key is then used to estimate the gain and quantum bit error rate (QBER).
(a) Key rates per pulse as a function of mean photon number ($\mu$) for two cases: (i) traditional analytical technique and (ii) proposed numerical technique using simulated results
(b) Simulated key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol.
(c) Experimental secure key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol. Lines drawn are for the aid of the eye.
(d) Key rates per pulse as a function of distance (in km) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol. The plots are in log scale
Figure 3: (a) shows a comparison of secure key rates obtained using the traditional analytical method and the proposed numerical method. (b),(c) and (d) shows the comparison of secure key rates of the four different protocols using proposed analytical method.
V Results and discussion
We compare the traditional Decoy state method with the proposed SDP-based method in calculating lower bounds on the secure key rates given in Eq  (11). This comparison is performed for various values of the mean photon number as shown in Figure. (2(a)). This comparison is done by performing a simulation of the decoy-based BB84 protocol for various mean photon numbers. In the simulation, the mean photon number of the decoy is kept constant with $\nu=0.1$. It is seen that our proposed method gives tighter lower bounds than the traditional analytic method by $27\%$ at the optimal mean photon number ($\mu=0.8$).
We employ our SDP-based method to analyse the key rates for the four protocols viz. BB84, decoy state protocol, coincidence detection (CD) protocol, and decoy state coincidence detection (DSCD) protocol. We use the SDP technique to find the bounds on the yields and errors to compute the bounds on key rates using Eqs. (11) and (5). The respective key rates are plotted as a function of the mean photon number ($\mu$) in Figure. (2(b)). It is established that when coincidences are taken into account, the key rate is enhanced by a substantial amount. The coincidence detection protocol helps to achieve a higher optimal mean photon number, i.e. $\mu=1.1$, as compared to using the traditional decoy-state method without coincidences, i.e. $\mu=0.8$. Comparing the optimal key rates of decoy-state protocol and DSCD protocol, we note an improvement of $\approx 64\%$.
Using the experimental setup described in the previous Section, the four protocols are executed. The protocols are systematically repeated for several signal intensities, each corresponding to distinct mean photon numbers. The experimentally acquired gain ($Q_{\mu}$) and QBER ($E_{\mu}$) are then used to estimate the optimal bounds on the secure key rate as discussed in Section. III. We have used the timestamps for recording the coincidences in addition to the single detections. The data were taken at various intensities of the source. These two cases enabled us to study four variants for our QKD protocol: Standard BB84, BB84 with decoy, BB84 with coincidences, BB84 with decoy and coincidences. Optimal key rates were calculated for all four cases at different values of mean photon number ($\mu$). The key rates for all the cases as a function of the mean photon number can be seen in Figure. (2(c)). It is also observed that our experiment reproduces the theoretical predictions of improvement in secure key rates. We obtained a maximum key rate of $0.17$ per pulse for the DSCD protocol.
We have performed a simulation of the expected secure key rate for the four protocols as a function of distance in optical fiber. The simulated key rates are plotted in Figure. (2(d)). In this simulation we have considered the mean photon number of the signal and decoy to be $\mu=0.9$ and $\nu=0.1$, respectively. It is evident that the DSCD protocol consistently gives higher key rates than other protocols across varying distances, showcasing its effectiveness in extending communication ranges. This consistent performance highlights the DSCD protocol’s reliability and strength in addressing challenges related to communication distance.
VI Conclusion
In this work, we have proposed the decoy state coincidence detection (DSCD) protocol, where we have integrated two approaches to harness their collective strength. Monitoring coincidences in decoy state protocol allows us to detect the most practical, sophisticated PNS attacks. With an assurance of no PNS attack, we showed, including contributions from two-photon statistics that we can achieve higher key rates.
A novel method for optimizing key rates using Semi-Definite Programming (SDP) is introduced, providing tight bounds on asymptotic key rates for the protocols discussed in this paper. In practical terms, this method yields reasonable bounds on key rates within a few minutes when executed on a personal computer. Moreover, the method is readily parallelizable, ensuring adaptability for scenarios where computed bounds may need further tightening, although we have not encountered such a situation in our experience thus far. This method has been used to compute the asymptotic key rates for the field implementation of our protocol. The obtained results illustrate that employing the proposed protocol substantially enhances the asymptotic key rates. We plan to conduct a future study on the finite-size analysis of our proposed protocol.
Acknowledgement
The authors acknowledge the partial funding support from DST through the QuST program. The authors are also thankful to group members of the QST lab for their valuable input, with a special thanks to Dr. Satyajeet Patil and Mr. Vardaan Mongia. RB acknowledges support from Quantum Communications Hub of the UK Engineering and Physical Sciences Research Council (EPSRC). RB thanks Mr. Lewis Wooltorton, Dr. Shashaank Ranu and Prof. Roger Coleck for insightful discussions. AB thanks Prof. Marco Lucamarini and Dr. Rupesh Kumar for insightful discussions.
Disclosure
The authors declare no conflicts of interest.
Appendix A Ideal protocol
In this section, we discuss the definitions of yields and error rates for completeness. Furthermore, following the work of Ma et al. Ma2005 we state the expected values of these yields and error rates in an "ideal protocol", in which no eavesdropping is performed, and errors arise naturally due to lossy channels.
In our protocol, it is assumed that Alice prepares the signal state using phase-randomised weak coherent pulses given by:
[…]
$=$ $\sum_{n=0}^{\infty}\frac{\mu^{n}}{n!}\mathrm{e}^{-\mu}|n\rangle\langle n|,$
where $|\mu e^{i\theta}\rangle$ is the coherent state of the electromagnetic field fox_quantum . It is evident from Eq. (13), that the photon number $N$ for the weak coherent pulses, follow the poison distribution with probability $\mathbb{P}(N=n)$ given by Eq.(14).
$\mathbb{P}(N=n)=\frac{\mu^{n}}{n!}\mathrm{e}^{-\mu}.$ (14)
We adopt identical notations and formalism as presented in Ma2005 for congruence. Let $\eta$ be the overall efficiency for detection by Bob of a single photon sent by Alice. In an ideal protocol, the efficiency of detection of a $k$ photon pulse is given as:
$\eta_{k}=1-(1-\eta)^{k}\qquad\text{for }k\in\mathbb{N}$ (15)
Yield: Yield $Y_{k}$ is the conditional probability of a detection event at Bob’s side. The yield $Y_{k}$ includes the probability of detection of signal photons $\eta_{k}$ and the background events $Y_{0}$. We assume that the background counts and the signal are independent of each other. Then for an ideal protocol,
$Y_{k}$ $=Y_{0}+\eta_{k}-Y_{0}\eta_{k}$ (16)
Gain: Gain $Q_{\mu}$ is the probability with which a signal sent by Alice is detected by Bob. It depends on the characteristics of the source, channel and detectors, and hence is a function of mean photon number $\mu$ and the yield $Y_{i}$. The overall gain for any protocol is given in terms of yields by:
$Q_{\mu}$ $=$ $\sum_{i=0}^{\infty}Y_{i}\frac{\mu^{i}}{i!}e^{-\mu}=\sum_{i=0}^{\infty}Q_{i}$ (17)
where,
$Q_{i}$ $=$ $Y_{i}\frac{\mu^{i}}{i!}e^{-\mu}$ (18)
Quantum Bit Error Rate (QBER): Quantum Bit error rate (QBER) $E_{\mu}$ is the rate with which Bob makes a wrong detection even when his basis is compatible with Alice’s. The error rate of an k-photon state $e_{k}$ for an ideal protocol is given as,
$e_{k}=\frac{e_{0}Y_{0}+e_{d}\eta_{k}}{Y_{k}}.$ (19)
For any protocol (not necessarily an ideal protocol), QBER is related to mean photon number $\mu$, yields $Y_{k}$ and error rates $e_{k}$ as
$E_{\mu}Q_{\mu}$ $=$ $\sum_{k}^{\infty}e_{k}Y_{k}\frac{\mu^{k}}{k!}e^{-\mu}$ (20)
Appendix B Analytic bounds for key-rates
For completeness, we discuss the traditional method used to compute analytical bounds on the decoy state protocol. The secure key generation rate for the decoy based protocol is given by the equation
$\begin{split}R&\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\mathrm{bin}}(E_{\mu})+Q_{1}\left(1-H_{\mathrm{bin}}(e_{1})\right)\Big{\}},\end{split}$ (21)
The constraints for the problem are identical to those in the optimization problem. The key rate of the protocol is a function of several parameters, including the yields of single photon state $Y_{1}$ (or equivalently, the gain $Q_{1}$) and the corresponding error rate $e_{1}$. However, these are not experimentally observable in the QKD protocol. What can be experimentally obtained is the overall gain $Q_{\mu}$ and the Quantum Bit Error Rate (QBER) $E_{\mu}$. The decoy state method achieves secure QKD by providing a method to establish rigorous bounds on the parameters $Q_{1}$ and $e_{1}$ and hence establish bounds on the secure key rate. Using the decoy state method the following bounds have been derived
Appendix C Numerically computing key rates
Here we present an algorithm to compute tight lower bounds for the optimization problem (6).
C.1 Converting into a finite optimization problem
As the objective function of problem (6) only consists of $2$ terms with $4$ variables, removing infinite parameters from constraints suffices. Consider the following lower bound on the constraint for $Q_{\nu_{d}}$ in the optimization problem:
$Q_{\nu_{d}}e^{\nu_{d}}$ $=\sum_{k=0}^{n}Y_{k}\frac{\nu_{d}^{k}}{k!}+\sum_{k=n+1}^{\infty}Y_{k}\frac{\nu_{d}^{k}}{k!}$ (22)
$\leq\sum_{k=0}^{n}Y_{k}\frac{\nu_{d}^{k}}{k!}+\sum_{k=n+1}^{\infty}\frac{\nu_{d}^{k}}{k!}$
where the above inequality holds because $Y_{k}\in[0,1]$. Therefore, defining $\Theta_{n}(\nu_{d})$ by the sum
$\Theta_{n}(\nu_{d}):=\sum_{k=n+1}^{\infty}\frac{\nu_{d}^{k}}{k!}=e^{\nu_{d}}-\sum_{k=0}^{n}\frac{\nu_{d}^{k}}{k!}$ (23)
allows us to obtain a relaxation corresponding to the constraint $Q_{\mu}e^{\mu}=\sum_{k=0}^{\infty}Y_{k}\mu^{k}/k!$ in optimization problem (6):
$Q_{\mu}e^{\mu}-\Theta_{n}(\mu)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\mu}e^{\mu}.$ (24)
In other words, we have formed a relaxation of the constraint given by an infinite sum of infinitely many variables through a constraint that involves a finite sum of finitely many variables. We use similar arguments to derive the following relaxations of the constraints for $E_{\mu}Q_{\mu},Q_{\nu_{i}},E_{\nu_{i}}Q_{\nu_{i}}$ in optimization problem (6):
$E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}-\Theta_{n}(\nu_{d})\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}\leq E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}$ (25)
The above relaxations allow us to prove the following lemma:
Lemma 1.
Let $S=\{\mu\}\cup\{\nu_{i}\}_{i=1}^{m}$, then consider the optimization problem
$r_{n}:=\inf$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$ (26)
s.t. $\forall i\in\{0,\cdots,n\}:\quad Y_{i},e_{i}\in[0,1]$
$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}$
$\forall\lambda\in S:\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}$
$\forall\lambda\in S:\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
Then $r_{n}\leq r$. Furthermore,
$\lim_{n\rightarrow\infty}|r_{n}-r|=0$
Proof.
Consider the optimization problems for $r_{n}$ and $r$. From the preceding discussion, it is evident that the constraints in problem (26) serve as relaxations for the constraints in problem (6). Consequently, the feasible set of optimization problem (6) is a subset of that of optimization problem (26). Moreover, the optimization problems designed to compute both $r$ and $r_{n}$ share identical objective functions. Thus, $r_{n}\leq r$ holds for all $n\in\mathbb{R}$.
Additionally, it is straightforward to recognize that when $n<m$, the constraints of the optimization problem for $r_{n}$ are relaxations of those for $r_{m}$. This immediately implies that $r_{n}\leq r_{m}$. As we continue this process, the feasible set for the optimization problem for $r_{n}$ approaches that of the optimization problem for $r$. Consequently, the non-decreasing sequence $\{r_{n}\}_{n}$ will converge to $r$. ∎
C.2 Partitioning the domain
In this section, we rigorously define a partition of a set $\mathcal{D}$ and then demonstrate how partitioning the parameter set of an optimization problem can be employed to provide lower bounds for any general optimization problem. We subsequently utilize this technique to lower bound asymptotic key rates in the next section.
Consider an optimization problem of the form:
$k:=\min$ $f(\mathbf{x},\mathbf{y})$ (27)
$\mathrm{s.t.}$ $\mathbf{x}\in\mathcal{D}$
$\mathbf{y}\in\mathcal{Q}$
$\forall i\in\{1,\cdots,N\}:g_{i}(\mathbf{x},\mathbf{y})\leq 0$
$\forall j\in\{1,\cdots,M\}:h_{j}(\mathbf{x},\mathbf{y})=0$
where $\mathcal{D}\subset\mathbb{R}^{n}$, $\mathcal{Q}\subset\mathbb{R}^{m}$ and $f,g_{1},\cdots g_{N},h_{1},\cdots h_{M}$ are smooth functions defined on the domain $\mathcal{D}\times\mathcal{Q}$. We seek to compute lower bounds on $k$.
We now formally define a partition of a set $\mathcal{D}\subset\mathbb{R}^{n}$.
Definition 1 (Partition).
Let $\mathbf{a}=(a_{1},a_{2},\cdots a_{n}),\mathbf{b}=(b_{1},b_{2},\cdots,b_{n})\in\mathbb{R}^{n}$ be two distinct points such that $\forall i:a_{i}\leq b_{i}$. We adopt the notation $\mathcal{C}_{\mathbf{a},\mathbf{b}}:=[a_{1},b_{1}]\times[a_{2},b_{2}]\cdots[a_{n},b_{n}]$, where $\mathcal{C}_{\mathbf{a},\mathbf{b}}$ represents a (hyper) rectangle with $\mathbf{a}$ and $\mathbf{b}$ as the opposite vertices. We refer to a collection of (hyper) rectangles $\mathcal{P}=\{\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}\}_{i}$ as a partition (or grid) of the set $\mathcal{D}\subset\mathbb{R}^{n}$ if $\bigcup_{i}\mathcal{C}_{\mathbf{a}_{i}\mathbf{b}_{i}}\supseteq\mathcal{D}$.
For a partition $\mathcal{P}$, the following result provides a reliable lower bound on $k$.
Lemma 2.
Let $k$ be defined via optimization problem (27). Let $\mathcal{P}=\{\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}\}_{i}$ be any partition of the set $\mathcal{D}$. For each $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$, let $\tilde{f}_{i}:\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}\times\mathcal{Q}\mapsto\mathbb{R}$ be any lower bound of $f$ in the domain $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$. Then corresponding to each (hyper) rectangle $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$, define a sub-optimization problem:
$k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}):=\min$ $\tilde{f}_{i}(\mathbf{x},\mathbf{y})$ (28)
$\mathrm{s.t.}$ $\mathbf{x}\in\mathcal{C}_{\mathbf{a}_{i},\mathbf{b}_{i}}$
$\mathbf{y}\in\mathcal{Q}$
$\forall i\in\{1,\cdots,N\}:g_{i}(\mathbf{x},\mathbf{y})\leq 0$
$\forall j\in\{1,\cdots,M\}:h_{j}(\mathbf{x},\mathbf{y})=0$
Then the following holds:
[…]
Furthermore, if $\underset{{\mathbf{x}\in\mathcal{D},\mathbf{y}\in\mathcal{Q}}}{\sup}|f-\tilde{f}|\leq\epsilon$, then $k-k_{\mathcal{P}}\leq\epsilon$.
Proof.
Let $\mathbf{x}^{*},\mathbf{y}^{*}$ be optimal for the problem (27). Then $\mathbf{x}^{*}\in\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$ for some $\mathbf{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$. Furthermore, we define $\mathbf{x}_{i},\mathbf{y}_{i}$ to be the optimal for optimization problem (28). Thus, we have that
$k_{\mathcal{P}}\leq k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})=\tilde{f}(\mathbf{x}_{i},\mathbf{y}_{i})\leq\tilde{f}_{i}(\mathbf{x}^{*},\mathbf{y}^{*})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})=k.$ (30)
Here the first inequality (starting from the left to right) follows from the definition of $k_{\mathcal{P}}$. The first equality and the second inequality follows from the definition of $\mathbf{x}_{i},\mathbf{y}_{i}$. The third inequality follows from the fact that $\tilde{f}$ is a lower bound on $f$ in $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$. The final equality follows from the definition of $\mathbf{x}^{*},\mathbf{y}^{*}$.
To prove the second part, define
$k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}):=\min$ $f_{i}(\mathbf{x},\mathbf{y})$ (31)
$\mathrm{s.t.}$ $\mathbf{x}\in\mathcal{C}_{\mathbf{a}_{i},\mathbf{b}_{i}}$
$\mathbf{y}\in\mathcal{Q}$
$\forall i\in\{1,\cdots,N\}:g_{i}(\mathbf{x},\mathbf{y})\leq 0$
$\forall j\in\{1,\cdots,M\}:h_{j}(\mathbf{x},\mathbf{y})=0$
Clearly $k=\min_{i}\{k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}$. However, for every cuboid $\mathcal{C}_{\mathbf{a}^{(j)},\mathbf{b}^{(j)}}$, we have that
$k_{1}(\mathcal{C}_{\mathbf{a}^{(j)},\mathbf{b}^{(j)}})-k(\mathcal{C}_{\mathbf{a}^{(j)},\mathbf{b}^{(j)}})\leq f(\mathbf{x}_{j},\mathbf{y}_{j})-\tilde{f}(\mathbf{x}_{j},\mathbf{y}_{j})\leq\epsilon.$
here $\mathbf{x}_{j},\mathbf{y}_{j}$ are such that $k_{1}(\mathcal{C}_{\mathbf{a}^{(j)},\mathbf{b}^{(j)}})=f(\mathbf{x}_{j},\mathbf{y}_{j})$.
This implies that
$k=\min_{i}\{k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})-\epsilon\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}-\epsilon\geq k_{\mathcal{P}}-\epsilon.$
C.3 Algorithm for computing key-rates
As $e_{1}$ and $e_{2}$ both vary in the interval $[0,1]$, we construct a partition $\mathcal{P}=\bigcup_{i}\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$ of the set $[0,1]\times[0,1]$.
For each cuboid $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$, we define constants $\lambda_{k,i}^{\max}$ as via:
$\xi^{\max}_{k,i}=\begin{cases}\max\{\Phi((2a_{k}^{(i)}-1)^{k}),\Phi((2b_{k}^{(i)}-1)^{k})\}&\text{if}\quad 0\notin[2a_{k}^{(i)}-1,2b_{k}^{(i)}-1]\\ \Phi(0)=1&\text{otherwise}\end{cases}.$ (32)
As we shall show below, these constants are defined such that $\xi^{\max}_{k,i}$ tightly upper bound $\Phi((2e_{k}-1)^{k})$ in the rectangle $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$.
Lemma 3.
Let $f(e_{1},e_{2},Y_{1},Y_{2}):=\sum_{k\in\{1,2\}}Y_{k}\mu^{k}/k!(1-\Phi((2e_{k}-1)^{k}))$. Then $f_{i}(Y_{1},Y_{2},e_{1},e_{2}):=\sum_{k\in\{1,2\}}Y_{k}\mu^{k}/k!(1-\xi^{\max}_{k,i})$ is a lower bound on $f(e_{1},e_{2},Y_{1},Y_{2})$ in the cuboid $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$.
Proof.
As $Y_{k}$ are non-negative, to show that $f_{i}(\mathbf{x},\mathbf{y})\leq f(\mathbf{x},\mathbf{y})$ in $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$, it suffices to show that $\lambda^{\max}_{k,i}$ is an upper-bound of $\Phi((2e_{k}-1)^{k})$ (and hence $1-\xi^{\max}_{k,i}$ is a lower-bound of $1-\Phi((2e_{k}-1)^{k})$).
This is done by noticing certain properties of $\Phi(x)$. In particular, it is useful to note that
  • $\Phi^{\prime}(x)<0$ if $x\in(0,1)$.
  • $\Phi^{\prime}(x)=-\Phi^{\prime}(x)$.
Using these properties we see that
$\Phi^{\prime}(x^{k}):=\frac{\mathrm{d}\Phi(x^{k})}{\mathrm{d}x}=kx^{k-1}\Phi^{\prime}(x^{k})$
Thus, $\Phi^{\prime}(x^{k})$ also has the same two aforementioned properties:
  • $\Phi^{\prime}(x^{k})$ if $x\in(0,1)$.
  • $\Phi^{\prime}((-x)^{k})=k(-x)^{k-1}\Phi^{\prime}((-x)^{k})=(-1)^{2k-1}kx^{k-1}\Phi^{\prime}(x^{k})=-\Phi^{\prime}(x^{k})$.
To bound $\Phi((2x-1)^{k})$, Suppose $12a_{k}^{(i)}-$ 1 and $12b_{k}^{(i)}-$ 1 both have the same sign (i.e. they are both positive or both negative), then $\Phi((e_{k}-1)^{k})\leq\xi^{\max}_{k,i}$ follows from the fact that $\Phi(x^{k})$ is a monotonic function in $x\in(-1,0)$ and $x\in(0,1)$. If $12a_{k}^{(i)}-$ 1 and $12b_{k}^{(i)}-$ 1 have different sign, then it means that $0\in[2a_{k}^{(i)}-1,2b_{k}^{(i)}-1]$. As $\Phi(0)=1$, and $\Phi(x^{k})\leq 1$ implies the bound $\Phi((e_{k}-1)^{k})\leq\xi^{\max}_{k,i}$ holds for this case as well. ∎
Using the results above, we are now in the position to prove the following result:
Theorem 1.
Let $\mathcal{P}=\{\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}\}_{i}$ be a partition of the set $[0,1]\times[0,1]$ into rectangles defined as in Appendix C.2. Let $\xi^{\max}_{1,i}$, and $\xi^{\max}_{2,i}$ be defined according to eqns (32) and $\Theta_{n}(\lambda)$ be any upper bound of the sum $\sum_{i=n+1}^{\infty}(\lambda^{i}/i!)$. Define the optimization problem $r_{n}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})$ be the polynomial optimization problem
$r_{n}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}):=\inf\quad$ $Y_{1}\mu\left(1-\xi^{\max}_{1,i}\right)+Y_{2}\mu^{2}/2\left(1-\xi^{\max}_{2,i}\right)$ (33)
$\mathrm{s.t.}\quad\quad$ $e_{1}\in[a_{1},b_{1}]$
$e_{2}\in[a_{2},b_{2}]$
$\forall i\in\{0,\cdots,n\}:\quad Y_{i}\in[0,1]$
$\forall i\in\{3,\cdots,n\}:\quad e_{i}\in[0,1]$
$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
then
[…]
Proof.
The proof begins by first lower-bounding $r$ using $r_{n}$, as established in Lemma 1. Subsequently, we apply Lemma 2 to compute lower bounds on $r_{n}$. Specifically, let $\mathbf{x}=(e_{1},e_{2})$ and $\mathbf{y}=(Y_{0},Y_{1},\cdots,Y_{n},e_{0},e_{2},\cdots,e_{n})$ denote the parameters. Thus, $\mathcal{D}=[0,1]\times[0,1]$ and $\mathcal{Q}=[0,1]\times[0,1]\times\cdots\times[0,1]$ must be chosen.
[…]
Now, let $\mathcal{P}$ be a partition such that each rectangle $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$ is a rectangle of dimensions $\frac{1}{4N_{1}}\times\frac{1}{4N_{2}}$ (in other words, we chose a grid with uniform spacing). Note that the grid is chosen such that either $1/2\notin(2a_{k}^{(i)}-1,2b_{k}^{(i)}-1)$. Therefore, $\xi^{\max}_{k,i}$ is either $12a_{k}^{(i)}-$ 1 or $12b_{k}^{(i)}-$ 1.
Now, let $\epsilon>0$. As $r_{n}\rightarrow r$, chose $n$ such that $|r_{n}-r|\leq\epsilon$. Using monotonicity properties of $\Phi(x)\in[2a_{k}^{(i)}-1,2b_{k}^{(i)}-1]$, we obtain $|\Phi(2x-1)-\Phi(2a_{k}^{(i)}-1)|\leq|\Phi(2b_{k}^{(i)}-1)-\Phi(2a_{k}^{(i)}-1)|$. Similarly $|\Phi(2x-1)-\Phi(2b_{k}^{(i)}-1)|\leq|\Phi(2b_{k}^{(i)}-1)-\Phi(2a_{k}^{(i)}-1)|$. Thus, in the rectangle $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$, we get that
$f-\tilde{f}\leq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}|\Phi((2b_{k}^{(i)}-1)^{k})-\Phi((2a_{k}^{(i)}-1)^{k})|\leq\chi\sum_{k\in\{1,2\}}|\Phi((2b_{k}^{(i)}-1)^{k})-\Phi((2a_{k}^{(i)}-1)^{k})|$ (35)
where $\chi=\max\{\mu,\mu^{2}/2\}$.
[…]
C.4 Solving polynomial optimization problems
The challenge now remains to solve the optimization problem (33). Notice that, by design, this optimization problem is a polynomial optimization problem. Such problems are not inherently convex optimization problems; however, they can be relaxed to obtain a converging hierarchy of semi-definite programs (SDPs). Interested readers can refer to the excellent slides by Prof. Hamza Fawzi Hamza along with Section II of the article Polynomial_optimization and find further references for SDP relaxations of polynomial optimization problems. The resulting SDPs can be solved via well-known efficient algorithms.
In this work, we use the NCPOL2SDPA software NCPOL2SDPA to obtain SDP relaxations of polynomial optimization problems. The software directly computes these relaxations given a polynomial optimization problem and then calls an SDP solver directly, outputting the result. The MOSEK software mosek was employed to solve the Semi-Definite Programs used for computing the key rates.
Appendix D Speeding up solving the optimization problem
Note that in the optimization problem in (33), it is not necessary to partition the entire domain $[0,1]\times[0,1]$. This is because not all combinations of $e_{1}$ and $e_{2}$ are feasible within the other imposed constraints. Indeed, there is a maximal value of $e_{1}$ that complies with all the constraints, which was a fundamental element for establishing bounds on the key rate in the seminal paper by Ma et al. Ma2005 . Consequently, it is sufficient to consider $e_{1}$ within the reduced range $[0,e_{1}^{\text{up}}]$, where the upper limit $e_{1}^{\text{up}}$ is determined by the following expression:
$e_{1}^{\mathrm{up}}:=\max\quad$ $e_{1}$ (36)
s.t. $\forall k\in\{0,\cdots,n\}:\quad Y_{k},e_{k}\in[0,1]$
$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
Similarly, we can compute the upper bound for $e_{2}$, denoted as $e_{2}^{\text{up}}$, by formulating and solving the corresponding optimization problem. This procedure allows us to confine the feasible region for $e_{2}$ within the constraints of the problem:
$e_{2}^{\mathrm{up}}:=\max\quad$ $e_{2}$ (37)
s.t. $\forall k\in\{0,\cdots,n\}:\quad Y_{i},e_{i}\in[0,1]$
$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
$e_{1}\in[0,e_{1}^{\mathrm{up}}]$
In practice, we find that $e_{1}^{\textup{up}}\ll 1$ and $e_{2}^{\textup{up}}\ll 1$ which allows us to save a significant amount of computation time, rendering this method more practical.
Appendix E Rates in the presence of statistical errors
When estimating key rates, it is crucial to account for the fact that gains and quantum bit error rates (QBERs) are calculated based on finite statistical data. Consequently, there exists a small, albeit non-zero, probability that the experimental estimates of these values may deviate from their true values. To address such uncertainties, a common approach involves introducing a security parameter known as the completeness error.
Formally, the completeness error is defined as the probability that a protocol aborts in an ideal implementation. Ideally, a small completeness error is preferred. The completeness error, in conjunction with the number of rounds in the QKD protocol, is used to establish confidence intervals for experimental statistics. Consequently, the reported rates are determined by the chosen value completeness error. Opting for a smaller completeness error enhances the security of the protocol but results in lower certified rates.
Before proceeding to the discussion on confidence intervals for the statistics on gain and QBER, we recall that expression (5) is only applicable if the coincidences monitored in the protocol align with the expected (theoretically computed) coincidence distribution. In a protocol with finitely many rounds, it becomes necessary to ensure that the observed (single and multi-photon coincidences) are within a confidence interval from the theoretically expected distribution. This parameter $\epsilon_{\mathrm{stat}}$is pre-declared and ideally should be set as small as possible. The protocol should abort if the observed (single and multi) photon coincidences do not align with the expected values within the confidence interval $\epsilon_{\mathrm{stat}}$. In this paper, we do not consider the complete analysis of the statistical analysis for coincidence detections and leave it for future work.
We now return to the discussion on addressing issues of statistical uncertainties in inferring gains and QBERs. For the sake of simplicity in this section, the signal state is labelled as $\mu=\nu_{0}$. This means that we will not differentiate between the signal state and the decoy state in our analysis for this section, as the approach is similar for both cases. Moreover, the number of decoy states is represented by $K$. Recall that the Alice and Bob generate keys from $\mathcal{A}_{i}\in\{0,1\}$ and $\mathcal{B}_{i}\in\{0,1,\perp\}$, where $i\in\mathfrak{h}$ (which is the set of rounds that are not discarded – i.e. $\mathfrak{h}=\{i\in\{1,2,\cdots,n\}:\mathcal{X}_{i}=\mathcal{Y}_{i}\}$). Finally, to keep track of which decoy (or signal ) was sent, we defined the random variable $\mathcal{D}_{i}\in\{0,1,\cdots,K\}$. Furthermore, define the following random variables
$\mathcal{W}_{i}^{\nu_{d}}=\begin{cases}1&\text{if }\mathcal{D}_{k}=\nu_{d},\mathcal{A}_{i}\neq\mathcal{B}_{i}\\ 0&\text{otherwise}\end{cases}$ (38)
$\mathcal{U}_{i}^{\nu_{d}}=\begin{cases}1&\text{if }\mathcal{D}_{i}=\nu_{d},\mathcal{B}_{i}\neq\perp\\ 0&\text{otherwise}\end{cases}$ (39)
The gains $Q_{\nu}$ are estimated as $Q_{\nu_{d}}^{\mathrm{obs}}$ and $E_{\nu}^{\mathrm{obs}}$ that can be computed as
$Q_{\nu_{d}}^{\mathrm{obs}}$ ::=:= |$\frac{|\{i:\mathcal{U}_{i}^{\nu_{d}}=1\}|}{N_{\nu_{d}}}$ (40)
$E_{\nu_{d}}^{\mathrm{obs}}$ ::=:= |$\frac{|\{i:\mathcal{W}_{i}^{\nu_{d}}=1\}|}{N_{\nu_{d}}}.$ (41)
Where |.||.|| . | is used to denote the cardinality of the sets and $N_{\nu_{d}}$ is the number of rounds for which signal/decoy state $\nu_{d}$ has sent and the round was not discarded (i.e. $\mathcal{X}_{i}=\mathcal{Y}_{i}$ was recorded). For our protocol, the abort condition is given by:
$Q_{\nu_{d}}^{\mathrm{obs}}\in[Q_{\nu_{d}}^{\mathrm{exp}}-\delta_{\nu_{d}}^{Q},Q_{\nu_{d}}^{\mathrm{exp}}+\delta_{\nu_{d}}^{Q}]$ (42)
$E_{\nu_{d}}^{\mathrm{obs}}\in[E_{\nu_{d}}^{\mathrm{exp}}-\delta_{\nu_{d}}^{E},E_{\nu_{d}}^{\mathrm{exp}}+\delta_{\nu_{d}}^{E}]$ (43)
where $Q_{\nu_{d}}^{\mathrm{exp}}$ and $E_{\nu_{d}}^{\mathrm{exp}}$ are the true values of gains and QBERs, and $\delta_{\nu_{d}}^{Q}$ and $\delta_{\nu_{d}}^{E}$ are the values of statistical tolerances that we are willing to admit in the experimentally observed gains and QBERs. The tolerances $\delta_{\nu_{d}}^{Q}$ and $\delta_{\nu_{id}^{Q}}$ should be determined using the completeness error. To do so, we use the Hoeffding’s inequality, which we state here for convenience:
Lemma 4 (Hoeffding’s inequality).
Let $X_{i}$ be $n$ i.i.d. random variables with $a\leq X_{i}\leq b$, $a,b\in\mathbb{R}$. If $S=\sum_{i}X_{i}$ and $\mu=\mathbb{E}(S)$. Then for $t>0$
$\mathbb{P}(|S-\mu|\geq t)\leq 2\operatorname{e}^{-\frac{2t^{2}}{n(b-a)^{2}}}\,.$ (44)
Here $\mathbb{P}$ and $\mathbb{E}$ are the symbols used to denote probability and expected value of events, respectively.
Theorem 2.
Suppose the honest protocol that behaves in an i.i.d. fashion for $n=\sum_{d=0}^{K}N_{\nu_{d}}$ rounds. Let $\{\delta_{\nu_{d}}^{Q}\}_{d=0}^{K}$ and $\{\delta_{\nu_{d}}^{E}\}_{d=0}^{K}$ be the accepted tolerances in gains and QBERs that define the abort conditions (42) and (43). The probability that an honest protocol aborts is no greater than
$2\sum_{d=0}^{K}\left(e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{Q})^{2}}+e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{E})^{2}}\right)+2\epsilon_{\mathrm{stat}}.$ (45)
Proof.
The probability that the honest implementation aborts is given by the probability of the event $\mathfrak{A}$ given by
$\mathfrak{A}:=\left(\bigcup_{d}\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\}\right)\cup\left(\bigcup_{d}\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\}\right)\cup\mathfrak{A}_{1}.$ (46)
Where $Q_{\nu_{d}}^{\mathrm{exp}}$ and $E_{\nu_{d}}^{\mathrm{exp}}$ are the true values of Gain and QBER and $\mathfrak{A}_{1}$ is the event when observed values for coincidences are outside the confidence interval from ideal value, despite the true values being within the confidence interval. This can happen almost with the probability $2\epsilon_{\mathrm{stat}}$.
We can get a bound on $\mathbb{P}(\mathfrak{A})$ using the union bound
$\mathbb{P}(\mathfrak{A})\leq\sum_{d}\left(\mathbb{P}(\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\})+\mathbb{P}(\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\})\right)+\mathbb{P}(\mathbb{A}_{1}).$ (47)
We can bound the above by using Hoeffding’s inequality by taking $S=\sum_{i}\mathcal{W}_{i}$ or $S=\sum_{i}\mathcal{U}_{i}$:
($\mathbb{P}(\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\})\leq 2e^{-2n(\delta_{\nu_{d}}^{Q})^{2}}$ (48)
($\mathbb{P}(\{|E_{\nu_{i}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\})\leq 2e^{-2n(\delta_{\nu_{d}}^{E})^{2}}.$ (49)
Therefore, the following holds:
$\mathbb{P}(\mathfrak{A})\leq 2\sum_{s=0}^{K}\left(e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{Q})^{2}}+e^{-2N_{\nu_{i}}(\delta_{\nu_{d}}^{E})^{2}}\right)+2\epsilon_{\mathrm{stat}}.$ (50)
Using the result above, we can get a relationship between the completeness error $\epsilon_{C}$ and $\delta_{\nu_{d}}^{Q}$ or $\delta_{\nu_{d}}^{E}$ by the setting
$\forall i:\quad e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{Q})^{2}}=e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{E})^{2}}=\frac{\epsilon_{C}-2\epsilon_{\mathrm{stat}}}{2(K+1)}.$ (51)
Therefore, we can set
$\delta_{\nu_{d}}^{Q}=\delta_{\nu_{d}}^{E}=\sqrt{\frac{-1}{2N_{\nu_{d}}}\ln\left(\frac{\epsilon_{C}-2\epsilon_{\mathrm{stat}}}{2(K+1)}\right)}.$ (52)
Using this, we revise the constraints 24 and 25 to the following constraints
$(Q_{\nu_{d}}-\delta_{\nu_{d}}^{Q})e^{\nu_{d}}-\Theta_{n}(\nu_{d})$$\leq$ $\sum_{k=0}^{n}Y_{k}\frac{\nu_{d}^{k}}{k!}\leq(Q_{\nu_{d}}+\delta_{\nu_{d}}^{Q})e^{\nu_{d}}$ (53)
$(E_{\nu_{d}}-\delta_{\nu_{d}}^{E})(Q_{\nu_{d}}-\delta_{\nu_{d}}^{Q})e^{\nu_{d}}-\Theta_{n}(\nu_{d})$$\leq$ $\sum_{k=0}^{n}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}\leq(E_{\nu_{d}}+\delta_{\nu_{i}}^{E})(Q_{\nu_{d}}+\delta_{\nu_{d}^{Q}})e^{\nu_{d}}.$ (54)
The constraints for Theorem 1 should be modified accordingly to compute key-rate in the presence of experimental uncertainties. Note that by doing so, we receive a lower key-rate, as (53) and (54) are relaxations of 24 and 25 respectively.

### v2

@@
 We improve on the conventional approach of using decoy pulses by integrating it with the coincidence detection (CD) protocol.
-Additionally, we introduce an easy-to-implement algorithm to compute asymptotic key rates for the protocol.
+Additionally, we employ an easy-to-implement algorithm to compute asymptotic key rates for the protocol.
 Furthermore, we have carried out an experimental implementation of the protocol, where we demonstrate that monitoring coincidences in the decoy state protocol leads to enhanced key rates under realistic experimental conditions.
@@
 An adversary can exploit the presence of multiple photons in a pulse to perform a photon number splitting (PNS) attack to gain information on the shared key.
-To address the exposed vulnerabilities in the practical implementation of QKD protocols, we can either transition to a more robust protocol or apply security patches for a known attack.
+To address the exposed vulnerabilities in the practical implementation of QKD protocols, we can either transition to a more robust protocol or apply security patches for known attacks.
 Security patching involves monitoring various parameters of our system and assessing potential information leaks.
@@
 The integration of decoy states and coincidence detections helps us to strategically leverage their capabilities to our advantage.
-We will call this decoy state coincidence detection (DSCD) protocol.
+We will refer to these extra pulses as entrapped ones to avoid confusion with the decoy state protocol.
+We will call this the entrapped pulse coincidence detection (EPCD) protocol.
+This integration of entrapped pulses and coincidence detections allows us to strategically exploit their functionalities.
 If the obtained coincidence distribution indicates that a PNS attack is not carried out, we incorporate the contributions from two-photon gain ($Q_{2}$) and error ($e_{2}$) into the asymptotic key rate(refer to Lim:17 for the derivation of key rates in terms of $Q_{2}$ and $e_{2}$).
@@
 II Protocol
-The implementation of the BB84 protocol is susceptible to PNS attacks, where Eve exploits occasional multi-photon pulses generated by Alice’s source.
+The implementation of the BB84 protocol with weak coherent pulses is susceptible to PNS attacks, where Eve occasionally exploits multi-photon pulses generated by Alice’s source.
 The conventional countermeasure to the PNS attack, namely the decoy state protocol, involves modifications at Alice’s end by sending occasional decoy pulses alongside signal pulses.
@@
 biswas_quantum_2022 .
-The protocol proposed in this paper implements modifications at both Alice’s and Bob’s ends.
-Decoy states are sent from Alice’s end, and Bob’s end is equipped to monitor coincidence detections, effectively addressing PNS attacks and significantly enhancing key rates.
+The protocol proposed in this paper implements modifications at both Alice’s and Bob’s ends, aiming for higher key rates.
+Entrapped pulses are sent from Alice’s end, and Bob’s end is equipped to monitor coincidence detections, effectively addressing PNS attacks and significantly enhancing key rates.
 For clarity, the protocol is systematically stated below:
-Protocol (Decoy State Coincidence Detection Protocol (DSCD)).
-Here, we outline a general decoy state protocol, including coincidences monitoring.
-Alice transmits a signal with mean photon number $\nu_{0}$ accompanied by $K$ random decoy pulses with mean photon numbers $\nu_{1},\nu_{2},\cdots\nu_{K}$.
-Given a sequence of $n$ pulses, each labelled with an index $i$, the following steps are repeated for each pulse.
+Protocol (Entrapped Pulse Coincidence Detection Protocol (EPCD)).
+In this protocol, we present a comprehensive approach to utilizing entrapped pulses, along with monitoring coincidences.
+Alice sends a signal characterized by an average photon number $\nu_{0}$, along with $K$ entrapped pulses characterized by average photon numbers $\nu_{1},\nu_{2},\cdots,\nu_{K}$.
+For a sequence of $n$ pulses, each identified by an index $i$, the subsequent actions are iterated for each pulse.
   1. 1.
-    State preparation (Alice’s lab): Alice generates random numbers $\mathcal{D}_{i}\in\{0,1,\cdots K\}$, $\mathcal{X}_{i}\in\{0,1\}$ and $\mathcal{A}_{i}\in\{0,1\}$. Where,
+    State preparation (Alice’s lab): Alice generates random numbers $\mathcal{D}_{i}\in\{0,1,\cdots K\}$, $\mathcal{X}_{i}\in\{0,1\}$ and $\mathcal{A}_{i}\in\{0,1\}$ (Fig. 2). Where,
     1. (a)
-      $\mathcal{D}_{i}$ determines whether a signal ($\nu_{0}$) or a decoy-state ($\nu_{1},\nu_{2},\cdots\nu_{K}$) is transmitted. If $\mathcal{D}=d$ then mean photon number is $\nu_{d}$
+      $\mathcal{D}_{i}$ determines whether a signal ($\nu_{0}$) or an entrapped pulse ($\nu_{1},\nu_{2},\cdots\nu_{K}$) is transmitted. If $\mathcal{D}=d$ then mean photon number is $\nu_{d}$
     2. (b)
@@
       $\mathcal{A}_{i}$ is the bit value encoded for pulse $i$.
+    Figure 2: Graphical representation of signal and entrapped pulses randomly transmitted by Alice
   2. 2.
@@
 At wing, $\mathfrak{B}^{(0)}$ ($\mathfrak{B}^{(1)}$) Bob performs measurement in the Standard (Hadamard) basis.
-Let $\mathcal{B}_{i}^{(j)}$ denote the measured bit value for the pulse $i$ in port $j$.
+In our experiment, Standard is the Horizontal/Vertical basis, and Hadamard is the Diagonal/Antidiagonal basis.
+Let $\mathcal{B}_{i}^{(j)}$ denote the measured bit value for the pulse $i$ in port $j$ (Fig.
+3).
+    Figure 3: Graphical representation of basis selection and recorded measurements at Bob’s end for $i^{th}$ pulse
   4. 4.
@@
     1. (a)
-      If $\mathcal{B}_{i}^{(j\oplus 1)}=\perp$ and $\mathcal{B}_{i}^{(j)}\in\{0,1\}$: $\mathcal{Y}_{i}=j$ , $\mathcal{B}_{i}=\mathcal{B}_{i}^{(j)}$.
+      If $\mathcal{B}_{i}^{(j\oplus 1)}=\perp$ and $\mathcal{B}_{i}^{(j)}\in\{0,1\}$: $\mathcal{Y}_{i}=j$ , $\mathcal{B}_{i}=\mathcal{B}_{i}^{(j)}$. This indicates a single detection.
     2. (b)
       If $\mathcal{B}_{i}^{(0)},\mathcal{B}_{i}^{(1)}\in\{0,1\}$, then Bob sets $\mathcal{Y}_{i}=x$ and $\mathcal{B}_{i}=\mathcal{B}_{i}^{(x)}$ if $\mathcal{X}_{i}=x$.
+This indicates two-fold coincidence detection.
     3. (c)
@@
   8. 8.
-    Alice also discloses the decoy states used in each round $(\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n})$.
+    Alice also discloses the entrapped pulses used in each round $(\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n})$.
   9. 9.
@@
     $Q_{\nu_{d}}$ ::=:= |$\frac{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d,\mathcal{B}_{i}\neq\perp\}|}{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d\}|}$ (1)
-    and QBERs (the probability that $\mathcal{A}_{i}\neq\mathcal{B}_{i}$ whenever $\mathcal{X}_{i}=\mathcal{Y}_{i}$ ):
+    and QBERs (the probability that $\mathcal{A}_{i}\neq\mathcal{B}_{i}$ whenever $\mathcal{X}_{i}=\mathcal{Y}_{i}$, i.e. an incorrect measurement for a correct choice of basis):
     $E_{\nu_{d}}$ ::=:= |$\frac{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d,\mathcal{A}_{i}\neq\mathcal{B}_{i}\}|}{|\{i\in\mathfrak{g}:\mathcal{D}_{i}=d\}|}$ (2)
@@
 However, if the coincidences detected show the expected statistics of coincidences, then we can safely assume that the PNS attack is not carried out, and the key rate of Protocol Protocol can be calculated using the following bound (given in Lim:17 ):
-$R\geq\frac{1}{2}\left\{-Q_{\mu}H_{\text{bin}}(E_{\mu})f(E_{\mu})+Q_{1}\left(1-\Phi\left(2e_{1}-1\right)\right)+Q_{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\right\}.$ (5)
+$\begin{split}R\geq\frac{1}{2}\biggl{\{}&-Q_{\mu}H_{\text{bin}}(E_{\mu})f(E_{\mu})\\ &+Q_{1}\left(1-\Phi\left(2e_{1}-1\right)\right)\\ &+Q_{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\biggr{\}}.\end{split}$ (5)
 Here, $Q_{1}=Y_{1}\mu$ and $Q_{2}$ and $e_{2}$ are the two-photon gains and error rates (see appendix A), respectively.
@@
 Decoy state statistics aid in obtaining better estimates for both signal and decoy states, resulting in higher key rates.
-The same argument also applies in our protocol, where having the decoy states, in principle, also allows for better estimates for the single and two-photon statistics $Q_{1},Q_{2},e_{1}$ and $e_{2}$.
-Note that, as pointed out in felix_faint_2001 ; biswas_quantum_2022 ; Lim:17 , the PNS attack itself can be detected by simply measuring the coincidences even in the absence of decoy states.
+The same argument also applies in our protocol, where having the entrapped pulses, in principle, also allows for better estimates for the single and two-photon statistics $Q_{1},Q_{2},e_{1}$ and $e_{2}$.
+Note that, as pointed out in felix_faint_2001 ; biswas_quantum_2022 ; Lim:17 , the PNS attack itself can be detected by simply measuring the coincidences even in the absence of entrapped pulses.
 However, there are modifications to the PNS attack that cannot be detected simply by monitoring coincidences.
 One such example is an attack in which Eve only allows for one photon to pass and then sends photons herself to match the ideal coincidence detection.
-However, such an attack is easily prevented by using decoy states.
-This is because the adversary does not fundamentally know whether the state sent is a decoy or not; therefore, she will not be able to match the coincidences correctly when Alice and Bob condition on the rounds in which the signal/decoy was sent.
+However, such an attack is easily prevented by using entrapped pulses in addition to monitoring coincidences.
+This is because the adversary cannot distinguish between the entrapped pulses and the signal state; therefore, she will not be able to match the coincidences correctly.
 However, the strongest form of the modifications in the PNS attack will still not be detected using our protocol.
 In this attack, Eve injects a photon only if she stores one.
-However, we argue that such an attack is not practical for Eve to implement given the current technology.
+However, we argue that such an attack is not practical for Eve to implement, given the current technology.
 For Eve to execute such an attack, she must first determine that she has stored a photon, generate a replacement photon, and then send it to Bob.
@@
 In other words, the key rate can be obtained by solving the following optimization problem:
-$r:=\min$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$ (6)
-$\mathrm{s.t.}$ $\forall k:\quad Y_{k},e_{k}\in[0,1]$
+$r:=\min\Big{(}$ $Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)$
+$+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$
+s.t. $\forall k:\quad Y_{k},e_{k}\in[0,1]$
 $\forall d:\quad Q_{\nu_{d}}e^{\nu_{d}}=\sum_{k=0}^{\infty}Y_{k}\frac{\nu_{d}^{k}}{k!}$
-$\forall d:\quad E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}=\sum_{k=0}^{\infty}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}$
+$\forall d:\quad E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}=\sum_{k=0}^{\infty}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}$ (6)
 The reason for performing such an optimization is straightforward to understand.
@@
 However, what is known are the overall gain $Q_{\mu}$ and the QBER ($E_{\mu}$).
-The constraints of the optimization problem are related to the observed Gain and QBER for both the signal and decoy states.
+The constraints of the optimization problem are related to the observed Gain and QBER for both the signal and entrapped pulses.
 The fundamental idea is that the experimentally observed QBER and Gain should originate from certain, albeit unknown, fixed values of Yields and error rates.
@@
 Ma2005 conducted the first comprehensive exploration of lower bounds on the asymptotic key rates for the decoy state protocol.
-For the decoy state protocol without any coincidence detection, the optimization problem for computing the key rate differs from (6) only in the objective function.
+For the decoy state protocol without any coincidence detection, the optimization problem for computing the key rate differs from (III) only in the objective function.
 Specifically, it does not include contributions from the two-photon terms $Y_{2}$ and $e_{2}$ and is given by:
@@
 $Y_{0}\leq Y_{0}^{U}$ ::=:= $\frac{E_{\nu}Q_{\nu}e^{\nu}}{e_{0}},$ (8)
-$Y_{1}\geq Y^{L}_{1}$ ::=:= $\frac{\mu}{\mu\nu-\nu^{2}}\left(Q_{\nu}e^{\nu}-Q_{\mu}e^{\mu}\left(\frac{\nu^{2}}{\mu^{2}}\right)-Y_{0}^{U}\left(\frac{\mu^{2}-\nu^{2}}{\mu^{2}}\right)\right),$ (9)
+$Y_{1}\geq Y^{L}_{1}$ ::=:= $\frac{\mu}{\mu\nu-\nu^{2}}\left(Q_{\nu}e^{\nu}-Q_{\mu}e^{\mu}\left(\frac{\nu^{2}}{\mu^{2}}\right)\right.$ (9)
+$\left.-Y_{0}^{U}\left(\frac{\mu^{2}-\nu^{2}}{\mu^{2}}\right)\right),$
 $e_{1}\leq e_{1}^{U}$ ::=:= $\frac{E_{\nu}Q_{\nu}e^{\nu}}{Y_{1}^{L}\nu}$ (10)
 Exploiting the monotonicity properties of binary entropy, specifically that $H_{\mathrm{bin}}(x)$ is increasing for all $x\in[0,1/2]$, and the fact that $(1-H_{\mathrm{bin}}(x))>0$, a lower bound on the secure key rate is given by:
-$R\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\mathrm{bin}}(E_{\mu})+Y_{1}^{L}\mu e^{-\mu}\left(1-H_{\mathrm{bin}}(e_{1}^{U})\right)\Big{\}}.$ (11)
+[…]
+$+Y_{1}^{L}\mu e^{-\mu}\left(1-H_{\mathrm{bin}}(e_{1}^{U})\right)\Big{\}}.$ (11)
 An immediate approach to computing lower bounds on the key rate for the protocol with coincidence detection involves employing similar analytic techniques to establish upper and lower bounds, denoted as $Y_{2}^{L}$ and $e_{2}^{U}$, on $Y_{2}$ and $e_{2}$ respectively, that are compatible with the constraints.
 However, this process is tedious, and tight bounds may not be easily found.
-While numerical methods can be employed to find such bounds (one approach to obtain such bounds is shown in Appendix D), the results obtained are not sufficiently tight, as demonstrated in (refer to Figure-2(a).
+While numerical methods can be employed to find such bounds (one approach to obtain such bounds is shown in Appendix D), the results obtained are not sufficiently tight, as demonstrated in (refer to Figure-5(a).
 In the next section, we introduce a simple-to-implement method to iteratively derive bounds on the key rate that converge to the asymptotic key rate from below.
 III.1 Solving the optimization problem
-Two primary challenges emerge when addressing optimization problem (6).
-First, the problem is non-convex due to the non-convex nature of both the constraints and the objective function in (6).
+Two primary challenges emerge when addressing optimization problem (III).
+First, the problem is non-convex due to the non-convex nature of both the constraints and the objective function in (III).
 General optimization problems are notoriously difficult to solve due to their abstract nature unless they fit specific classes of convex optimization problems.
 As a result, the optimization problem cannot be solved directly using conventional techniques.
-Second, optimization problem (6) involves infinitely many free parameters, namely, $Y_{k}$ and $e_{k}$.
+Second, optimization problem (III) involves infinitely many free parameters, namely, $Y_{k}$ and $e_{k}$.
 The presence of these infinitely many parameters adds further complexity to the problem.
 The challenge of dealing with infinitely many variables can be addressed straightforwardly by recognizing that the contribution of $Y_{k}$ and $e_{k}$ when $k\gg 1$ is negligible to the sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ and $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}e_{k}$.
-Consequently, we relax the constraints of (6) by truncating the infinite sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ to involve sums over a finite number of variables.
+Consequently, we relax the constraints of (III) by truncating the infinite sums $\sum_{k}(\nu_{d}^{k}/k!)Y_{k}$ to involve sums over a finite number of variables.
 This comes, however, with a small penalty that depends on the number of terms kept in the sum.
@@
 We then use these tight bounds to lower-bound the objective function as follows
-$\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}(1-\Phi((2e_{k}-1)^{k}))\geq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}\cdot(1-\xi^{\max}_{k,i}).$ (12)
+${}\!\!\!\!\!\!\!\!\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}(1-\Phi((2e_{k}-1)^{k}))$
+$\geq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}\cdot(1-\xi^{\max}_{k,i}).$ (12)
 The loss in tightness when bounding the objective functions by such constants $\xi^{\max}_{1,i}$ and $\xi^{\max}_{2,i}$ depend upon the dimensions of the domain $\mathcal{C}_{i}$ as shown in Lemma 3, Appendix C.3.
@@
 However, making a finer partition comes at the expense of higher computation time, as more and more sub-problems need to be numerically solved.
-All the results lead to the following final result which forms the basis of the algorithm for computing key-rate $r$ (as defined in equation (6)). We state it here informally as:
+All the results lead to the following final result, which forms the basis of the algorithm for computing key rate $r$ (as defined in equation (III)). We state it here informally as:
 […]
@@
 The computation of finite-round statistics can, in principle, be performed using theoretical tools such as the Generalized Entropy Accumulation Theorem GEAT ; GEAT-security .
-Such analysis would require a more sophisticated analysis and is reserved for future work.
+Such analysis would require more sophisticated analysis and is reserved for future work.
 IV Experimental Method
@@
 (a) Key rates per pulse as a function of mean photon number ($\mu$) for two cases: (i) traditional analytical technique and (ii) proposed numerical technique using simulated results
-(b) Simulated key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol.
-(c) Experimental secure key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol.
+(b) Simulated key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) EPCD protocol.
+(c) Experimental secure key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) EPCD protocol.
 Lines drawn are for the aid of the eye.
-(d) Key rates per pulse as a function of distance (in km) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol. The plots are in log scale
+(d) Key rates per pulse as a function of distance (in km) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) EPCD protocol. The plots are in log scale
 Figure 5: (a) shows a comparison of secure key rates obtained using the traditional analytical method and the proposed numerical method.
-(b),(c) and (d) shows the comparison of secure key rates of the four different protocols using proposed analytical method.
+(b),(c) and (d) show the comparison of secure key rates of the four different protocols using the proposed analytical method.
 V Results and discussion
-We compare the traditional Decoy state method with the proposed SDP-based method in calculating lower bounds on the secure key rates given in Eq  (11).
+Increasing the key rate in realistic scenarios remains a challenge as alternative approaches typically demand additional hardware components or intricate optimization techniques for key extraction.
+We proposed a solution minimizing the resource requirement on both the software and hardware side.
+We performed a comparative study of key rates across different protocols.
+We then evaluated our key rate optimization method against existing ones, with detailed results presented in subsequent paragraphs.
+We compare the traditional Decoy state method with the proposed SDP-based method in calculating lower bounds on the secure key rates given in Eq  (III).
 This comparison is performed for various values of the mean photon number as shown in Figure.
-(2(a)).
-This comparison is done by performing a simulation of the decoy-based BB84 protocol for various mean photon numbers.
+(5(a)), which is done by performing a simulation of the decoy-based BB84 protocol for various mean photon numbers.
 In the simulation, the mean photon number of the decoy is kept constant with $\nu=0.1$.
@@
 We employ our SDP-based method to analyse the key rates for the four protocols viz.
-BB84, decoy state protocol, coincidence detection (CD) protocol, and decoy state coincidence detection (DSCD) protocol.
+BB84, decoy state protocol, coincidence detection (CD) protocol, and entrapped pulse coincidence detection (EPCD) protocol.
 We use the SDP technique to find the bounds on the yields and errors to compute the bounds on key rates using Eqs.
-(11) and (5).
+(III) and (5).
 The respective key rates are plotted as a function of the mean photon number ($\mu$) in Figure.
-(2(b)).
+(5(b)).
 It is established that when coincidences are taken into account, the key rate is enhanced by a substantial amount.
@@
 $\mu=0.8$.
-Comparing the optimal key rates of decoy-state protocol and DSCD protocol, we note an improvement of $\approx 64\%$.
+Comparing the optimal key rates of decoy-state protocol and EPCD protocol, we note an improvement of $\approx 64\%$ corresponding to higher achievable distances.
 Using the experimental setup described in the previous Section, the four protocols are executed.
@@
 The data were taken at various intensities of the source.
-These two cases enabled us to study four variants for our QKD protocol: Standard BB84, BB84 with decoy, BB84 with coincidences, BB84 with decoy and coincidences.
+These two cases enabled us to study four variants for our QKD protocol: Standard BB84, decoy state protocol, CD protocol, and EPCD protocol.
 Optimal key rates were calculated for all four cases at different values of mean photon number ($\mu$).
 The key rates for all the cases as a function of the mean photon number can be seen in Figure.
-(2(c)).
+(5(c)).
 It is also observed that our experiment reproduces the theoretical predictions of improvement in secure key rates.
-We obtained a maximum key rate of $0.17$ per pulse for the DSCD protocol.
+We obtained a maximum key rate of $0.17$ per pulse for the EPCD protocol.
 We have performed a simulation of the expected secure key rate for the four protocols as a function of distance in optical fiber.
 The simulated key rates are plotted in Figure.
-(2(d)).
+(5(d)).
 In this simulation we have considered the mean photon number of the signal and decoy to be $\mu=0.9$ and $\nu=0.1$, respectively.
-It is evident that the DSCD protocol consistently gives higher key rates than other protocols across varying distances, showcasing its effectiveness in extending communication ranges.
-This consistent performance highlights the DSCD protocol’s reliability and strength in addressing challenges related to communication distance.
+It is evident that the EPCD protocol consistently gives higher key rates than other protocols across varying distances, showcasing its effectiveness in extending communication ranges.
+This consistent performance highlights the EPCD protocol’s reliability and strength in addressing challenges related to communication distance.
 VI Conclusion
-In this work, we have proposed the decoy state coincidence detection (DSCD) protocol, where we have integrated two approaches to harness their collective strength.
-Monitoring coincidences in decoy state protocol allows us to detect the most practical, sophisticated PNS attacks.
-With an assurance of no PNS attack, we showed, including contributions from two-photon statistics that we can achieve higher key rates.
+In this work, we have proposed the entrapped pulse coincidence detection (EPCD) protocol, where we have integrated two approaches to harness their collective strength.
+Monitoring coincidences of the signal as well as entrapped pulses allows us to detect the most practical, sophisticated PNS attacks.
+With an assurance of no PNS attack, we showed that including contributions from two-photon statistics, that we could achieve higher key rates.
 A novel method for optimizing key rates using Semi-Definite Programming (SDP) is introduced, providing tight bounds on asymptotic key rates for the protocols discussed in this paper.
@@
 Lewis Wooltorton, Dr.
-Shashaank Ranu and Prof.
+Shashank Ranu and Prof.
 Roger Coleck for insightful discussions.
@@
 For completeness, we discuss the traditional method used to compute analytical bounds on the decoy state protocol. The secure key generation rate for the decoy based protocol is given by the equation
-$\begin{split}R&\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\mathrm{bin}}(E_{\mu})+Q_{1}\left(1-H_{\mathrm{bin}}(e_{1})\right)\Big{\}},\end{split}$ (21)
+$R$ $\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\text{bin}}(E_{\mu})$
+$\qquad+Q_{1}\left(1-H_{\text{bin}}(e_{1})\right)\Big{\}},$ (21)
 The constraints for the problem are identical to those in the optimization problem.
@@
 Appendix C Numerically computing key rates
-Here we present an algorithm to compute tight lower bounds for the optimization problem (6).
+Here we present an algorithm to compute tight lower bounds for the optimization problem (III).
 C.1 Converting into a finite optimization problem
-As the objective function of problem (6) only consists of $2$ terms with $4$ variables, removing infinite parameters from constraints suffices.
+As the objective function of problem (III) only consists of $2$ terms with $4$ variables, removing infinite parameters from constraints suffices.
 Consider the following lower bound on the constraint for $Q_{\nu_{d}}$ in the optimization problem:
@@
 $\Theta_{n}(\nu_{d}):=\sum_{k=n+1}^{\infty}\frac{\nu_{d}^{k}}{k!}=e^{\nu_{d}}-\sum_{k=0}^{n}\frac{\nu_{d}^{k}}{k!}$ (23)
-allows us to obtain a relaxation corresponding to the constraint $Q_{\mu}e^{\mu}=\sum_{k=0}^{\infty}Y_{k}\mu^{k}/k!$ in optimization problem (6):
+allows us to obtain a relaxation corresponding to the constraint $Q_{\mu}e^{\mu}=\sum_{k=0}^{\infty}Y_{k}\mu^{k}/k!$ in optimization problem (III):
 $Q_{\mu}e^{\mu}-\Theta_{n}(\mu)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\mu}e^{\mu}.$ (24)
 In other words, we have formed a relaxation of the constraint given by an infinite sum of infinitely many variables through a constraint that involves a finite sum of finitely many variables.
-We use similar arguments to derive the following relaxations of the constraints for $E_{\mu}Q_{\mu},Q_{\nu_{i}},E_{\nu_{i}}Q_{\nu_{i}}$ in optimization problem (6):
+We use similar arguments to derive the following relaxations of the constraints for $E_{\mu}Q_{\mu},Q_{\nu_{i}},E_{\nu_{i}}Q_{\nu_{i}}$ in optimization problem (III):
 $E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}-\Theta_{n}(\nu_{d})\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\nu_{d}^{k}}{k!}\leq E_{\nu_{d}}Q_{\nu_{d}}e^{\nu_{d}}$ (25)
@@
 Let $S=\{\mu\}\cup\{\nu_{i}\}_{i=1}^{m}$, then consider the optimization problem
-$r_{n}:=\inf$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$ (26)
+$r_{n}:=\inf$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)$
+$+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$
 s.t. $\forall i\in\{0,\cdots,n\}:\quad Y_{i},e_{i}\in[0,1]$
@@
 $\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}$
-$\forall\lambda\in S:\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
+$\forall\lambda\in S:\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$ (26)
 Then $r_{n}\leq r$. Furthermore,
@@
 Consider the optimization problems for $r_{n}$ and $r$.
-From the preceding discussion, it is evident that the constraints in problem (26) serve as relaxations for the constraints in problem (6).
-Consequently, the feasible set of optimization problem (6) is a subset of that of optimization problem (26).
+From the preceding discussion, it is evident that the constraints in problem (1) serve as relaxations for the constraints in problem (III).
+Consequently, the feasible set of optimization problem (III) is a subset of that of optimization problem (1).
 Moreover, the optimization problems designed to compute both $r$ and $r_{n}$ share identical objective functions.
@@
 Thus, we have that
-$k_{\mathcal{P}}\leq k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})=\tilde{f}(\mathbf{x}_{i},\mathbf{y}_{i})\leq\tilde{f}_{i}(\mathbf{x}^{*},\mathbf{y}^{*})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})=k.$ (30)
+$k_{\mathcal{P}}\leq k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})$ $=\tilde{f}(\mathbf{x}_{i},\mathbf{y}_{i})\leq\tilde{f}_{i}(\mathbf{x}^{*},\mathbf{y}^{*})$
+$\leq f(\mathbf{x}^{*},\mathbf{y}^{*})=k.$ (30)
 Here the first inequality (starting from the left to right) follows from the definition of $k_{\mathcal{P}}$.
@@
 To prove the second part, define
-$k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}):=\min$ $f_{i}(\mathbf{x},\mathbf{y})$ (31)
+$k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})$ $:=\min f_{i}(\mathbf{x},\mathbf{y})$
 $\mathrm{s.t.}$ $\mathbf{x}\in\mathcal{C}_{\mathbf{a}_{i},\mathbf{b}_{i}}$
@@
 $\forall i\in\{1,\cdots,N\}:g_{i}(\mathbf{x},\mathbf{y})\leq 0$
-$\forall j\in\{1,\cdots,M\}:h_{j}(\mathbf{x},\mathbf{y})=0$
+$\forall j\in\{1,\cdots,M\}:h_{j}(\mathbf{x},\mathbf{y})=0$ (31)
 Clearly $k=\min_{i}\{k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}$. However, for every cuboid $\mathcal{C}_{\mathbf{a}^{(j)},\mathbf{b}^{(j)}}$, we have that
@@
 This implies that
-$k=\min_{i}\{k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})-\epsilon\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}-\epsilon\geq k_{\mathcal{P}}-\epsilon.$
+$k$ $=\min_{i}\{k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})-\epsilon\}$
+$\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}-\epsilon\geq k_{\mathcal{P}}-\epsilon.$ (32)
 C.3 Algorithm for computing key-rates
@@
 For each cuboid $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$, we define constants $\lambda_{k,i}^{\max}$ as via:
-$\xi^{\max}_{k,i}=\begin{cases}\max\{\Phi((2a_{k}^{(i)}-1)^{k}),\Phi((2b_{k}^{(i)}-1)^{k})\}&\text{if}\quad 0\notin[2a_{k}^{(i)}-1,2b_{k}^{(i)}-1]\\ \Phi(0)=1&\text{otherwise}\end{cases}.$ (32)
+$\xi^{\max}_{k,i}=\begin{cases}\begin{aligned} &\max\{\Phi((2a_{k}^{(i)}-1)^{k}),\\ &\Phi((2b_{k}^{(i)}-1)^{k})\}\end{aligned}&\text{if}\quad 0\notin[2a_{k}^{(i)}-1,2b_{k}^{(i)}-1]\\ \Phi(0)=1&\text{otherwise}\end{cases}$ (33)
 As we shall show below, these constants are defined such that $\xi^{\max}_{k,i}$ tightly upper bound $\Phi((2e_{k}-1)^{k})$ in the rectangle $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$.
@@
 Thus, $\Phi^{\prime}(x^{k})$ also has the same two aforementioned properties:
-  • $\Phi^{\prime}(x^{k})$ if $x\in(0,1)$.
+  • $\Phi^{\prime}(x^{k})\leq 0$ if $x\in(0,1)$.
   • $\Phi^{\prime}((-x)^{k})=k(-x)^{k-1}\Phi^{\prime}((-x)^{k})=(-1)^{2k-1}kx^{k-1}\Phi^{\prime}(x^{k})=-\Phi^{\prime}(x^{k})$.
@@
 $r_{n}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})$ $:=\inf\quad Y_{1}\mu\left(1-\xi^{\max}_{1,i}\right)+Y_{2}\mu^{2}/2\left(1-\xi^{\max}_{2,i}\right)$ (34)
-$\mathrm{s.t.}\quad\quad$ $e_{1}\in[a_{1},b_{1}]$
+$\mathrm{s.t.}$ $e_{1}\in[a_{1},b_{1}]$
 $e_{2}\in[a_{2},b_{2}]$
@@
 $\forall i\in\{3,\cdots,n\}:\quad e_{i}\in[0,1]$
-$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
-$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
+$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:$
+$\quad\quad Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}$
+$\quad\quad\leq Q_{\lambda}\exp(\lambda)$
+$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:$
+$\quad\quad E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
+$\quad\quad\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
 then
@@
 Thus, in the rectangle $\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}$, we get that
-$f-\tilde{f}\leq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}|\Phi((2b_{k}^{(i)}-1)^{k})-\Phi((2a_{k}^{(i)}-1)^{k})|\leq\chi\sum_{k\in\{1,2\}}|\Phi((2b_{k}^{(i)}-1)^{k})-\Phi((2a_{k}^{(i)}-1)^{k})|$ (35)
+$f-\tilde{f}$ $\leq\sum_{k\in\{1,2\}}Y_{k}\frac{\mu^{k}}{k!}|\Phi((2b_{k}^{(i)}-1)^{k})-\Phi((2a_{k}^{(i)}-1)^{k})|$
+$\leq\chi\sum_{k\in\{1,2\}}|\Phi((2b_{k}^{(i)}-1)^{k})-\Phi((2a_{k}^{(i)}-1)^{k})|$
 where $\chi=\max\{\mu,\mu^{2}/2\}$.
@@
 s.t. $\forall k\in\{0,\cdots,n\}:\quad Y_{k},e_{k}\in[0,1]$
-$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
-$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
+$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
+$\quad\quad\quad\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
+$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
+$\quad\quad\quad\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
 Similarly, we can compute the upper bound for $e_{2}$, denoted as $e_{2}^{\text{up}}$, by formulating and solving the corresponding optimization problem.
@@
 s.t. $\forall k\in\{0,\cdots,n\}:\quad Y_{i},e_{i}\in[0,1]$
-$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
-$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
+$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
+$\quad\quad\quad\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
+$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
+$\quad\quad\quad\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
 $e_{1}\in[0,e_{1}^{\mathrm{up}}]$
@@
 The probability that the honest implementation aborts is given by the probability of the event $\mathfrak{A}$ given by
-$\mathfrak{A}:=\left(\bigcup_{d}\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\}\right)\cup\left(\bigcup_{d}\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\}\right)\cup\mathfrak{A}_{1}.$ (46)
+$\mathfrak{A}$ ::=:= $\left(\bigcup_{d}\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\}\right)$ (46)
+$\cup\left(\bigcup_{d}\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\}\right)\cup\mathfrak{A}_{1}.$
 Where $Q_{\nu_{d}}^{\mathrm{exp}}$ and $E_{\nu_{d}}^{\mathrm{exp}}$ are the true values of Gain and QBER and $\mathfrak{A}_{1}$ is the event when observed values for coincidences are outside the confidence interval from ideal value, despite the true values being within the confidence interval.
@@
 We can get a bound on $\mathbb{P}(\mathfrak{A})$ using the union bound
-$\mathbb{P}(\mathfrak{A})\leq\sum_{d}\left(\mathbb{P}(\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\})+\mathbb{P}(\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\})\right)+\mathbb{P}(\mathbb{A}_{1}).$ (47)
+$\mathbb{P}(\mathfrak{A})$ $\leq\sum_{d}\Bigl{(}\mathbb{P}(\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\})$
++($\quad+\mathbb{P}(\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\})\Bigr{)}+\mathbb{P}(\mathbb{A}_{1}).$
 We can bound the above by using Hoeffding’s inequality by taking $S=\sum_{i}\mathcal{W}_{i}$ or $S=\sum_{i}\mathcal{U}_{i}$:
@@
 Therefore, the following holds:
-$\mathbb{P}(\mathfrak{A})\leq 2\sum_{s=0}^{K}\left(e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{Q})^{2}}+e^{-2N_{\nu_{i}}(\delta_{\nu_{d}}^{E})^{2}}\right)+2\epsilon_{\mathrm{stat}}.$ (50)
+$\mathbb{P}(\mathfrak{A})$$\leq$ $2\sum_{s=0}^{K}\left(e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{Q})^{2}}+e^{-2N_{\nu_{i}}(\delta_{\nu_{d}}^{E})^{2}}\right)$ (49)
+$+2\epsilon_{\mathrm{stat}}.$
 Using the result above, we can get a relationship between the completeness error $\epsilon_{C}$ and $\delta_{\nu_{d}}^{Q}$ or $\delta_{\nu_{d}}^{E}$ by the setting